OpenStudy (anonymous):

(6/7)(-5/18)+120/7 divided by 45/3

OpenStudy (james1769):

u lookin for a way to solve or just a answer?

OpenStudy (anonymous):

How to solve

OpenStudy (james1769):

just break it down in to parts

OpenStudy (anonymous):

so 6/7 * -5/18 and you get -51/21 ?

OpenStudy (anonymous):

-51/21 + 120/7 = 103/7 ?

OpenStudy (anonymous):

103/7 divided by 45/3 + 103/105?

OpenStudy (anonymous):

Does the original equation look like this? \[\frac{ 6 }{ 7 }\times \frac{ -5 }{ 18 }+\frac{ 120 }{ 7 }\div \frac{ 45 }{ 3 }?\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

with no parenthesis?

OpenStudy (anonymous):

there are () but they mean the same as multiplying.

OpenStudy (anonymous):

So the only parenthesis are around 6/7 and -5/18?

OpenStudy (anonymous):

If that's the case, according to order of operations, what needs to happen first?

OpenStudy (anonymous):

You were right with multiplying \[\frac{ 6 }{ 7 }\times \frac{ -5 }{ 18 }\]

OpenStudy (anonymous):

you need to multiply (6/7)(-5/18) first i know this.

OpenStudy (anonymous):

right, and you know that when you multiply fractions, if possible, you can also simply it before multiplying (just to work with smaller numbers)

OpenStudy (anonymous):

do i have to find a common denominator?

OpenStudy (anonymous):

i know the answer is 19/21 but i do not understand how i got the answer

OpenStudy (anonymous):